### v1

# TimeSeries2Report prompting enables adaptive large language model management of lithium-ion batteries

Abstract
Large language models (LLMs) offer promising capabilities for interpreting multivariate time-series data, yet their application to real-world battery energy storage system (BESS) operation and maintenance remains largely unexplored. Here, we present TimeSeries2Report (TS2R), a prompting framework that converts raw lithium-ion battery operational time-series into structured, semantically enriched reports, enabling LLMs to reason, predict, and make decisions in BESS management scenarios. TS2R encodes short-term temporal dynamics into natural language through a combination of segmentation, semantic abstraction, and rule-based interpretation, effectively bridging low-level sensor signals with high-level contextual insights. We benchmark TS2R across both lab-scale and real-world datasets, evaluating report quality and downstream task performance in anomaly detection, state-of-charge prediction, and charging/discharging management. Compared with vision-, embedding-, and text-based prompting baselines, report-based prompting via TS2R consistently improves LLM performance in terms of across accuracy, robustness, and explainability metrics. Notably, TS2R-integrated LLMs achieve expert-level decision quality and predictive consistency without retraining or architecture modification, establishing a practical path for adaptive, LLM-driven battery intelligence.
1 Introduction
From portable electronics [liang2019review] to electric vehicles [cano2018batteries, harper2019recycling, liu2025multi] and grid-scale renewable energy storage systems [arteaga2017overview, jiang2025battery], lithium-ion batteries (LIBs) power nearly every facet of modern life. As deployment scales from individual cells to gigawatt-hour battery energy storage systems (BESS), ensuring the safety, reliability, and efficiency of such large, heterogeneous networks has become increasingly complex and mission-critical. Effective battery management requires precise, real-time monitoring of cell states including state of charge (SOC) [adaikkappan2022modeling, shen2017co], state of health (SOH) [hu2020battery, jones2022impedance, severson2019data], and temperature [cao2025model, mei2023operando] under highly dynamic load conditions, evolving degradation pathways, and environmental variability. Yet the coupled electrochemical, thermal, and mechanical dynamics across thousands of interacting cells make predictive maintenance and adaptive control fundamentally challenging [mama2025comprehensive, che2022state, xie2022faults].
Artificial intelligence (AI) has emerged as a powerful enabler of advanced battery management, supporting a wide range of tasks from diagnostics to optimization. Deep learning models such as long short-term memory (LSTM) networks [tan2019transfer], convolutional neural networks (CNNs) [gu2023novel, mazzi2024lithium], and physics-informed neural networks (PINNs) [wang2024physics, nascimento2021hybrid] have been successfully applied to estimate SOH and forecast battery aging. Unsupervised methods including autoencoders [cao2025model, zhang2023realistic], Gaussian mixture models [yang2025toward], and isolation forests [jiang2022fault] have proven effective for anomaly detection, identifying faults such as overcharging, short circuits, and connection failures by extracting latent patterns from operational data. Reinforcement learning approaches [wang2025data, park2022deep, hao2023adaptive] have further enabled dynamic energy and charging strategy optimization under varying load conditions.
Despite these advances, most AI-based battery management systems remain task-specific, data-hungry, and difficult to generalize across chemistries, architectures, or usage scenarios. Their limited interpretability and retraining overhead hinder deployment in safety-critical applications where transparency and adaptability are essential. These limitations are particularly pronounced in real-world operation and maintenance (OM) settings, where regulatory compliance, reliability, and human oversight are non-negotiable. While explainable AI (XAI) techniques such as SHapley Additive exPlanations (SHAP) [lee2022state, zhao2025lithium] and Layer-wise Relevance Propagation (LRP) [wang2023explainability] have been proposed to enhance model interpretability, they typically yield post hoc numerical feature attributions that require significant cognitive effort to interpret – falling short of true human-AI collaboration.
Large language models (LLMs) [naveed2025comprehensive, li2025self, dagdelen2024structured, guo2025deepseek, achiam2023gpt, zhao2025align, kevian2024capabilities] offer a compelling alternative, enabling flexible reasoning, contextual understanding, and natural language generation across diverse tasks. Their general-purpose nature opens the door to a new paradigm in battery management – one centered on explainability, adaptability, and task generalization. However, mainstream LLMs such as ChatGPT [achiam2023gpt] are designed primarily for text-based conversations and lack native support for interpreting continuous, multivariate, and temporally structured time-series data. As highlighted in recent studies [jiang2024empowering, wang2025itformer, 11268252], this architectural mismatch limits their utility in time-series-dominant domains such as battery OM, where key tasks like SOH estimation and fault diagnosis depend on rich temporal dynamics. Recent developments in multimodal LLMs, such as ChatTS [xie2024chatts] and Time-LLM [jin2023time], address this gap by learning shared embedding spaces for time-series and text, allowing LLMs to treat signal segments as pseudo-tokens. However, these models require extensive retraining and deep natural language processing (NLP) expertise, which limits their accessibility and practical use in the battery domain.
In this work, we introduce TimeSeries2Report (TS2R), an automated signal-to-semantics-to-report framework that enables LLMs to perform explainable battery management without specialized architecture or retraining. Our method converts raw operational time-series data such as voltage, current, temperature, and capacity into structured, human-readable reports that serve as textual proxies for the original signals. These reports are constructed using a universal template that generalizes across multivariate and multi-cell LIB systems, and can be directly integrated into off-the-shelf LLMs via prompt engineering. By translating quantitative dynamics into semantically rich descriptions, TimeSeries2Report empowers LLMs to carry out key OM tasks including condition assessment, anomaly detection, and decision support with strong interpretability and minimal infrastructure cost. Notably, our approach requires no model training, making it lightweight, scalable, and accessible to users beyond the AI community. To support reproducibility and further research, we release the first large-scale paired dataset of LIB time-series and descriptive reports, consisting of 1,792 time-series samples, each comprising 259,778 minutes, aligned with 72,520 reports, along with a benchmark suite covering three representative battery management tasks. All code and data are open-sourced.
2 Results
2.1 Illustration of the TimeSeries2Report framework
As illustrated in the top panel of Fig. 1a, a typical BESS consists of numerous LIB cells connected in series and/or parallel configurations. The first stage of the TS2R framework involves data acquisition and pre-processing. For each cell, we record key operational measurements such as voltage, current, and temperature and compute system-level statistics across all cells at each time. These aggregated metrics, including the mean, maximum, minimum, standard deviation, and entropy, characterize the aggregate behavior and heterogeneity of the BESS (bottom panel of Fig. 1a, Methods 4.1).
Subsequently, the time-series data from all individual cells and the corresponding system-level statistics are aggregated into a unified multivariate dataset. Each time series is then segmented into fine-grained, temporally ordered slices to capture short-term dynamical behavior. The segmentation window is chosen to be sufficiently short to preserve local trends while retaining meaningful temporal structure rather than noise [kil1997optimum]. In this work, the temporal window width is fixed at 10 time stamps, determined through expert knowledge and representative observations (see Methods 4.2). To characterize these short-term dynamics, we design a set of attributes that comprehensively describe local temporal behavior (see Methods 4.3 and Table 1). For each attribute of each slice, semantic descriptors are automatically assigned using a rule-based computational pipeline (see Methods 4.4 and Table 2). For example, the attribute “trend” is mapped to semantic descriptors such as increasing, stable, or decreasing. This procedure transforms raw numerical time series into sequences of semantic descriptors. Adjacent slices with identical semantic descriptors are then merged to reduce redundancy and enhance temporal coherence. This merging is performed across the full time span and all variables, yielding compact yet information-rich representations that retain the salient dynamical features of the system (see Methods 4.5). Finally, the consolidated descriptor sequences are translated into human-readable expressions using predefined phrasing rules (see Methods 4.6). At this stage (Fig. 1b), each time series is converted into a concise textual summary that integrates multiple dynamic attributes, thereby bridging quantitative time-series data with interpretable natural-language representations. Next, the framework synthesizes the generated expressions into a comprehensive report using an LLM (Fig. 1c). The LLM receives two inputs: (i) the textual expressions generated in the previous stage and (ii) detailed instructions specifying the report’s desired descriptive style and output structure (see Methods 4.7). By integrating these inputs, the LLM produces a coherent, context-aware report that describes the original time-series data from multiple dynamic-attribute perspectives.
The generated report serves as readily usable inputs for downstream LLM-driven OM tasks. We demonstrate their applicability in three representative scenarios: SOC prediction, operational monitoring (abnormality detection), and charging/discharging management (Fig. 1d and Methods 4.8). Given these sentences and the task-specific user queries, the LLM generates interpretable responses supported by explicit, semantically grounded reasoning, thereby enhancing transparency and explainability throughout the battery OM workflow.
2.2 Benchmarking the performance of TimeSeries2Report on open-source datasets of standalone LIB cells
To evaluate the performance of TS2R, we sourced two public datasets from Refs. [severson2019data] and [zhu2022data], and curated them into the MIT dataset and TJU dataset, respectively (see Supplementary Note 1, Table S1 and Fig. S1 for detailed curation steps). Both datasets consist of measurements from single, standalone LIB cells but differ in cell chemistry: the MIT dataset corresponds to lithium iron phosphate (LFP) cells, while the TJU dataset contains nickel cobalt aluminum oxide (NCA) cells. Since these datasets involve standalone cells rather than interconnected multi-cell systems, the computation of system-level statistics and downstream multi-cell operations is bypassed. Instead, we focus on evaluating the quality of reports generated for individual LIB cells across a range of operating conditions, including constant-current (CC) charging, CC-to-constant-voltage (CV) charging, idle-to-CC discharging, and CC discharging-to-idle transitions, as well as across chemistries (LFP and NCA) and LLM backbones (Qwen3-14B [yang2025qwen3], Llama3.1-8B [dubey2024llama], DeepSeek-v3.2 [guo2025deepseek], and ChatGLM-4.5 [zeng2025glm]).
In Fig. 2a, we present representative outputs based on the voltage time series of a single LIB cell, where the voltage increases continuously until the 55 time stamp and remains stable thereafter (top panel). The three columns compare: (i) the direct output of Qwen3‑14B prompted with raw numerical time-series data (left), (ii) the output from TS2R-processed input using the same LLM (middle, TS2R+Qwen3‑14B), and (iii) an expert-written reference report serving as ground truth (right). Unlike the raw-input baseline, TS2R+Qwen3‑14B accurately identifies both the global trend (initial increase followed by stability) and the transition point at the 55 time stamp. It also more consistently aligns with expert annotations across other descriptive attributes. In contrast, the baseline Qwen3-14B fails to capture these patterns reliably.
To quantify the generated report quality, we introduce FactScore [min2023factscore], a metric that assesses the alignment between generated reports and expert-written references (see Supplementary Note 2 for annotation protocol). As illustrated in Fig. 2b, each sentence generated for different signals (e.g., voltage, temperature, current) and attributes (e.g., trend, transition, outlier) is compared to its expert counterpart. An independent judging LLM (Qwen3-235B-A22B-Instruct-2507 [yang2025qwen3]) evaluates each sentence pair and assigns a score based on consistency: 1 for high consistency, 0.5 for partial consistency, and 0 for low consistency. The final FactScore is computed as the average score across all attribute-level evaluations for a given time series. Full judging prompts are detailed in Supplementary Note 3.
We quantitatively assess the performance of TS2R using FactScore on both the MIT (Fig. 2c) and TJU (Fig. 2d) datasets, covering a range of operating conditions and LLM backbones, including Qwen3‑14B, Llama3.1‑8B, DeepSeek‑v3.2 and ChatGLM-4.5. To evaluate effectiveness, we compare two prompting strategies: (i) directly inputting raw numerical time-series data as prompts into the LLM, and (ii) first converting the data into structured natural language using TS2R prior to prompting. Across all conditions, the TS2R-augmented approach consistently outperforms direct numerical prompting, yielding FactScore improvements of 6–57, most with statistical significance (). These results highlight the inherent difficulty text-based LLMs face when interpreting raw time-series data, particularly in capturing complex temporal features such as trends, transitions, and outliers. By translating numerical signals into interpretable natural language, TS2R effectively unlocks the semantic reasoning capabilities of general-purpose LLMs.
The consistent gains observed across all four models underscore the model-agnostic nature of TS2R, demonstrating that its benefits generalize across architectures without requiring model-specific tuning. Importantly, TS2R achieves this performance without any retraining, offering a lightweight, scalable, and adaptable solution for generating descriptive reports from single-cell time-series data. These findings position TS2R as a practical and interpretable framework for real-world battery management applications.
2.3 Evaluating TS2R in real-world, system-level settings
While the MIT and TJU datasets were collected under well-controlled laboratory conditions and focus exclusively on standalone LIB cells, evaluating TS2R in a realistic, system-level context requires more representative field data. To this end, we collected operational data from a deployed BESS consisting of 28 modules, each containing 16 LFP cells connected in series. Over a six-month period (December 1, 2023 to May 30, 2024), we recorded minute-level time-series measurements of voltage, temperature, and SOC from each individual cell, as well as module-level current. Since all cells within a module are connected in series, the module current is equivalent to the current flowing through each cell. In total, this yielded 28 modules × 16 cells × 4 variables 1,792 time series, each comprising 259,778 minutes. The dataset includes over 400 operational charge-discharge cycles, executed under constant-power (CP) protocols in response to dynamic grid demands. Each sample in this dataset, referred to as the ZJU‑o dataset (with “o” denoting omni), consists of synchronized time-series data for voltage, temperature, SOC, and current, capturing real-world, module-level battery behavior under diverse operating conditions (see Supplementary Note 4, Table S2 and Fig. S2 for data acquisition details). The difference in operational complexity between lab-scale and real-world deployments is illustrated in Fig. 3a and Fig. S3, which compare the distributions of charging and discharging currents in the MIT and ZJU‑o datasets. The ZJU‑o dataset exhibits substantially broader current variability, reflecting increased operational diversity and environmental uncertainty – factors that pose a more rigorous test for evaluating the robustness of TS2R-generated reports. Due to the sheer scale of the ZJU‑o dataset, manual annotation by human experts is impractical. To enable benchmark evaluation, we curated a smaller test set referred to as the ZJU‑t dataset (with “t” for test). This was constructed by: (i) randomly sampling a module index , (ii) selecting a random initial time within the operational timeline, and (iii) recording all four variables (voltage, temperature, SOC, current) for all 16 cells in module over the window minutes. This process was repeated 100 times, resulting in 100 time-series samples with expert-written annotations generated using the same protocol described in Supplementary Note 2. The ZJU‑t dataset captures module-level behavior by including multiple cells per module, allowing us to compute system-level statistics (note: while technically “module-level,” we use “system-level” hereafter without ambiguity). As shown in Fig. 3b, we aggregate per-cell measurements at each time stamp to compute summary statistics (e.g., mean and standard deviation of voltage), yielding system-level features that reflect both aggregate trends and inter-cell variability. This enables TS2R to operate hierarchically, interpreting battery behavior at both the individual cell and system level, thereby supporting hierarchical interpretation of battery system behavior.
In Fig. 3c, we present a representative report generated from the system-level voltage statistics shown in Fig. 3b. In this example, the average voltage exhibits a transition from a stable increase to a sharp decrease around the 6163 time stamp. When prompted with raw numerical data, the LLM (Qwen3‑14B) fails to capture this transition and incorrectly reports a monotonic increase over the entire period. In contrast, the same LLM, when provided with TS2R-parsed input (TS2R+Qwen3‑14B), correctly identifies the increasing phase, the sharp decrease, and the subsequent stabilization. The resulting report closely matches the expert-written reference in both qualitative trend description and quantitative accuracy. A similar effect is observed in Fig. 3d, where the LLM is prompted with the standard deviation of voltage: TS2R again enables the model to detect and articulate key statistical transitions that are overlooked when using raw numerical input alone.
To systematically assess performance, we applied the same evaluation protocol used in Fig. 2b to the ZJU‑t dataset, measuring report quality at both the single-cell and module (system) levels. Fig. 3e and Fig. S4a summarize the FactScore results for module-level reports, evaluated across nine descriptive attributes shown in Table 1, three operational states (charging, discharging, and idle), and four LLM backbones (Qwen3‑14B, Llama3.1‑8B, DeepSeek‑v3.2 and ChatGLM-4.5). Corresponding FactScore results for individual cells are provided in Supplementary Fig. S4b. Across all conditions, LLMs augmented with TS2R consistently outperform those using raw numerical input, except in cases where the baseline performance is already near saturation (for example, the fluctuation attribute in Fig. S4a), demonstrating significantly improved accuracy in capturing module-level dynamics. Among the tested models, both TS2R+DeepSeek‑v3.2 and TS2R+ChatGLM-4.5 achieve the remarkably promising performance, with FactScore metrics approaching 0.81.0 in most cases. These results confirm that TS2R substantially enhances LLMs’ ability to interpret complex, real-world time-series data and generate reports that align closely with expert-written references, both at the single-cell and system levels. This underscores the potential of TS2R to advance downstream battery OM tasks through accurate, interpretable, and scalable report generation.
2.4 TS2R enhances the accuracy of SOC prediction
We next demonstrate how TS2R can be integrated into and improve the performance of a critical downstream task in LIB management: SOC prediction, which is essential for understanding the charging and discharging status.
To ensure a fair comparison, we focus exclusively on LLM-based approaches, as outlined in Fig. 4a. The methods are grouped into four categories: (i) Text-based LLMs (Qwen3‑14B, ChatGLM‑4.5), which take raw numerical time-series values as direct textual input. (ii) Vision-based LLMs (Qwen3‑VL‑Plus [yang2025qwen3], ChatGLM‑4.5V [hong2025glm]), which interpret time-series data rendered as graphical plots. (iii) Time-series embedding-based models (ChatTS [xie2024chatts]), which encode numerical time series into dense vector representations aligned with language tokens. (iv) Report-based LLMs (TS2R+Qwen3‑14B, TS2R+ChatGLM‑4.5), which leverage TS2R to convert raw time-series data into structured natural-language reports. DeepSeek-v3.2 and Llama3.1-8B are excluded from the following analyses, as they do not have vision-based counterparts.
To evaluate the performance of the seven models across the four categories, we followed the assessment pipeline shown in Fig. 4b. Using the ZJU-t dataset (see Fig. 3e), we assessed prediction accuracy using two standard metrics: root mean squared error (RMSE) and mean absolute error (MAE). For each test sample, the first 90 minutes of the time series were provided as input, and each method was tasked with predicting the system-level average SOC at minute 100. RMSE and MAE were then calculated by comparing the predicted SOC with the ground truth at that time point. Prompting configurations for each method are detailed in Supplementary Note 5.
Example outputs from the seven SOC prediction methods are shown in Fig. S5. In Fig. 4c, we compare their prediction accuracy across the four LLM-based categories. Among these, the time-series embedding model (ChatTS) exhibits the lowest accuracy. This is likely due to its training on a highly curated dataset composed primarily of synthetically generated sequences, which limits its generalization to real-world data such as the ZJU-t dataset that features complex and diverse latent temporal patterns. Although retraining ChatTS on ZJU-t could improve its performance, doing so would require substantial computational resources and time. This limitation highlights a key advantage of TS2R: it enhances LLM performance without the need for retraining. Vision-based LLMs (Qwen3‑VL‑Plus and ChatGLM‑4.5V) rank second lowest in performance. While transforming numerical time series into visual plots helps preserve coarse trends, the resulting pixel-based representations introduce quantization artifacts that constrain the model’s ability to perform fine-grained reasoning, thereby reducing prediction precision. Text-based LLMs (Qwen3‑14B and ChatGLM‑4.5) perform moderately well, producing acceptable SOC predictions when directly prompted with raw numerical input. In contrast, report-based LLMs, which utilize TS2R-generated natural language descriptions, consistently achieve the highest prediction accuracy. Notably, the TS2R+Qwen3‑14B model reduces RMSE by 20 and 74 compared to its corresponding text-based and vision-based variants (Qwen3‑14B and Qwen3‑VL‑Plus, respectively). Similar improvements are observed with the ChatGLM-based models. These results underscore the effectiveness of TS2R in enabling LLMs to better interpret complex time-series data and perform accurate forecasting without the need for architecture modification or retraining.
2.5 TS2R enables accurate detection of abnormalities in LIB operation monitoring
A fundamental task in LIB operation monitoring is the real-time detection of abnormalities, enabling timely interventions to prevent safety hazards and performance degradation. The goal of anomaly detection is twofold: (i) to accurately identify true abnormal events, and (ii) to minimize false positives, i.e., to avoid misclassifying normal behavior as abnormal. We evaluated this task using the same seven LLM-based methods described earlier. Each method was first prompted to classify a given time series as either normal or abnormal. If an abnormality was detected, the model was further queried to specify the anomaly type (e.g., voltage-related, temperature-related). Prompting details are provided in Supplementary Note 5. Since the original ZJU-t dataset contains only normal operating conditions, it is insufficient for this evaluation. To address this, we constructed the ZJU-ta dataset (“a” indicating the inclusion of abnormal cases) by replacing the final three normal-operation samples in ZJU-t with three curated abnormal scenarios: (1) temperature abnormality in a single cell, (2) voltage abnormality in a single cell, and (3) voltage abnormality at the system level. Details of these abnormal cases are provided in Table S4. Importantly, the ratio of abnormal to normal samples reflects the low incidence of anomalies observed in real-world deployments. The evaluation pipeline is illustrated in Fig. 5a. For each sample, all methods were prompted to perform anomaly detection. Predictions were compared against expert annotations (ground truth), and two metrics were computed over the ZJU-ta dataset: Accuracy (Acc), measuring the ability to correctly identify true anomalies (objective i), and False Alarm Rate (FAR), quantifying the frequency of incorrectly labeling normal behavior as abnormal (objective ii).
Example outputs for both abnormal and normal samples across the seven LLM-based methods are presented in Figs. S6 and S7. Fig. 5b summarizes the quantitative comparison using two evaluation metrics Acc and FAR. Among all methods, the time-series embedding-based model (ChatTS) achieves an accuracy below 0.1 and FAR around 0.9, clearly inadequate for practical LIB operation monitoring. Interestingly, we observe significant performance differences even between models with the same LLM backbone (e.g., Qwen3-series vs. ChatGLM-series), likely due to architectural differences inherent to the base models. However, since such model-specific factors are not the focus of this study, we compare methods within each LLM family. For the Qwen3-series, the vision-based method (Qwen3-VL-Plus) achieves the lowest accuracy (0.03), while the report-based method (TS2R+Qwen3-14B) improves accuracy by more than a factor of two over the text-based variant (Qwen3-14B), and reduces FAR by 28.92. Similar trends are observed within the ChatGLM series. Notably, TS2R+ChatGLM-4.5 achieves the best overall performance, with an accuracy of approximately 0.9 and FAR below 0.1, making it the most suitable choice for real-world LIB operation monitoring. These findings are further supported by the qualitative comparison of model outputs for the three abnormal samples, detailed in Table S5.
2.6 TS2R supports more precise predictive decision-making for LIB charging and discharging management
Since TS2R has demonstrated strong predictive capabilities, we next investigate whether these benefits can be translated into improved decision-making. A key function of BESS management is to make timely decisions to terminate charging or discharging in order to prevent over-stressing or under-stressing individual cells – actions critical for both safe operation and long-term longevity of the system. Typically, BESS managing system uses the system-level maximum or minimum SOC, i.e., the highest or lowest SOC across all LIB cells, as the criterion to determine safe charging/discharging cutoffs. To illustrate how TS2R supports predictive decision-making for charging and discharging management, we apply two filtering conditions to the ZJU-t dataset: (i) the system must be mid-way through a charging cycle at minute 100, and (ii) the corresponding ZJU-o dataset must show that the system-level maximum SOC continues to rise after minute 100, eventually reaching 97, a cutoff defined by the BESS manufacturer. This second condition ensures that discharging events (e.g., grid demand response) do not interfere, allowing the true remaining charging time (RCT) to be recorded as ground truth for evaluating predictive accuracy. Using these constraints, we identify a representative sample (Sample 47) from the ZJU-t dataset, shown in Fig. 6a. In this sample, the system-level maximum SOC at minute 100 is 23 (marked by a blue star in Fig. 6a). The task is to predict the remaining charging time : how long it will take for the maximum SOC to reach the cutoff value of 97, given only time-series data up to minute 100. By referencing the corresponding ZJU-o record, the ground truth RCT is minutes (marked by a red star in Fig. 6a). Prediction performance is quantified by the relative error . A smaller indicates a more accurate predictive decision by the LLM-based method. Details of the prompting setup are provided in Supplementary Note 5. Fig. S8 presents sample predictions of RCT from the seven LLM-based methods.
Because LLMs are inherently generative and exhibit stochastic outputs, we repeated the relative error computation five times for each LLM-based method, following the evaluation protocol outlined in Fig. 6a. In Fig. 6b, we compare the resulting relative errors for RCT predictions across the seven methods. The embedding-based model ChatTS yields substantially higher values than either report-based method (TS2R+Qwen3‑14B and TS2R+ChatGLM‑4.5). Within each LLM family, the report-based variant consistently achieves the lowest relative error and tighter confidence intervals, indicating superior prediction accuracy and consistency. This demonstrates TS2R’s ability to enhance both the precision and reliability of LLM-based predictive decision-making. Notably, the combination of TS2R and ChatGLM‑4.5 delivers the most accurate and consistent predictions overall, with a remarkably low and narrow confidence band. These results underscore its strong potential for real-world deployment in BESS management systems.
3 Discussion
In summary, this work presents TimeSeries2Report, a prompting framework that enables LLMs to effectively interpret and manage BESS by transforming raw multivariate time-series data into structured, semantically enriched natural language reports. By bridging low-level operational measurements with high-level contextual understanding, TS2R enhances the interpretability and decision-making capability of general-purpose LLMs. Through extensive benchmarking on both lab-scale (MIT, TJU) and real-world datasets (ZJU-t), we demonstrate that TS2R substantially improves the factual consistency of generated reports across diverse BESS operating conditions (CC charging, CC charging-to-CV charging, idle-to-CC discharging, and CC discharging-to-idle) and across multiple LLM backbones (Qwen3-14B, Llama3.1-8B, ChatGLM-4.5, DeepSeek-v3.2), outperforming baselines that rely on raw numerical input.
Among all tested combinations, TS2R+ChatGLM-4.5 and TS2R+DeepSeek-v3.2 achieve the highest factual accuracy and consistency. Notably, given the overwhelming volume of BESS operational data, expert annotation at scale is prohibitively time-consuming, creating a major bottleneck in developing domain-specific machine learning models. The high-quality outputs generated by TS2R+ChatGLM-4.5 and TS2R+DeepSeek-v3.2 enable automatic annotation of large-scale BESS data, providing a viable path toward scalable dataset construction. Leveraging this capability and accounting for API-call costs, we use TS2R+DeepSeek-v3.2 to annotate the entire ZJU-o dataset, making it, to the best of our knowledge, the first large-scale, paired BESS operational dataset with natural language annotations. We release this resource to support future research in LLM-driven BESS monitoring, prediction, and autonomous decision-making.
We further demonstrate that TS2R’s ability to generate reports with high factual accuracy and consistency directly supports improved performance across a wide range of downstream OM tasks, including SOC prediction, anomaly detection, and charging/discharging management. In all these tasks, TS2R outperforms competing approaches that represent time-series data as plots, embeddings, or plain numerical text. These findings reinforce TS2R’s core motivation: enhancing LLM comprehension of complex data by translating it into natural language. Among all evaluated combinations, TS2R+ChatGLM-4.5 consistently achieves the best trade-off between predictive accuracy, interpretability, and robustness across tasks. This highlights its strong potential for real-world deployment in intelligent, adaptive, and trustworthy BESS management systems.
One of the most notable advantages of TS2R is its ability to enhance LLMs’ capacity to interpret complex time-series data without requiring retraining, fine-tuning, or architectural modifications – processes that are often computationally expensive and time-consuming. Because TS2R operates purely as a prompting framework, it can leverage off-the-shelf LLMs via standard API calls, offering a lightweight, scalable, and interpretable solution for battery OM that is readily deployable across diverse systems. While TS2R already achieves strong performance in zero-shot settings, pairing it with fine-tuned LLMs could further improve results by internalizing domain-specific knowledge for battery time-series reporting. In such a supervised fine-tuning setting, the LLM is updated using available time-series data and corresponding offline reports. With the time-series input, human instruction and semantic descriptors serving as prompts, the model is trained to align its output with the reference report. The fine-tuning procedure is described as follows. Let denote the input multivariate battery time-series data, where represents the number of time stamps, and is the number of variables at each time stamp. First, TS2R pipeline converts time-series data into semantic descriptors:
where is the generated descriptors, denotes the conversion pipeline. Then TS2R takes the human instruction (denoted as ) and the semantic descriptors as the input of LLM, and generates the corresponding time-series report :
where represents the function of LLM, and denotes its network parameters as the target of fine-tuning. The training objective follows a standard cross-entropy loss at the token level:
where is the length of the LLM’s generated report in tokens, is the vocabulary size, represents the one-hot encoded target probability for the token at the position in the report, denotes the predicted probability of the token at the position, as produced by the LLM.
For more advanced applications, TS2R can be integrated into a retrieval-augmented generation (RAG) framework to support long-context and domain-enriched decision-making. In this setup, generated reports are segmented (e.g., system-level vs. cell-level content), indexed, and stored in a vector database. When a user issues a query, the most relevant report segments are retrieved and passed to the LLM as external knowledge. This enables the LLM to formulate accurate, context-aware responses, especially valuable in large-scale LIB systems where a single report may be too long for one-shot processing. Additionally, domain knowledge (e.g., safety thresholds, manufacturer specs) can be seamlessly integrated into the RAG database to further support professional-grade OM tasks.
4 Methods
4.1 Derivation of system-level statistical time series
The first step of the TS2R framework is to compute system-level statistical time series from individual LIB cell measurements. At each time point, we calculate a set of statistics, including mean, maximum, minimum, standard deviation, and Shannon entropy, for every measured variable across all cells in the system. In the MIT dataset, these variables include voltage, current, temperature, and charge/discharge capacities; in the TJU dataset, they include voltage, current, charge capacity, and discharge capacity; and in the ZJU-series dataset, they include voltage, temperature, SOC, and current. These statistics characterize both the aggregate behavior and the heterogeneity of the battery system, providing a compact representation of its spatiotemporal dynamics. The resulting system-level features, together with the raw time-series data from individual cells, constitute the preprocessed inputs for the subsequent stages of the TS2R pipeline.
4.2 Partitioning time series into slices
Before semantic descriptors can be assigned, the raw time series must be decomposed into units that capture local, phase-specific behavior. Assigning a single descriptor to an entire sequence is typically inadequate, as LIB signals often exhibit complex, multi-stage dynamics. For example, a single time series may include multiple charging, discharging, and idle phases, making global labels such as increasing, decreasing, or stable insufficient to characterize the full trajectory. To address this, we partition each time series into multiple equal-length time slices. The slice length is chosen to be sufficiently short to preserve local dynamical patterns, ensuring that each segment reflects a relatively coherent behavior that can be uniquely and accurately characterized by the semantic attributes introduced in the next step.
4.3 Design of semantic attributes
To translate time-series data into natural language, we define a set of semantic attributes that specify the key informational elements to be captured in the generated report. These attributes determine which aspects of the time series should be described, such as trends, transitions, variability, and extremal values, and serve as the foundational units for producing interpretable and consistent expressions. The same semantic attribute framework is applied uniformly to both single-cell time series and system-level statistical time series, ensuring consistency across hierarchical levels of LIB monitoring.
To characterize both the behavior of individual cells (e.g., voltage, temperature) and the aggregate behavior of the entire system (mean, maximum, minimum for each variable across cells), we organize the semantic attributes into two categories: shape attributes and statistical attributes. Shape attributes capture temporal dynamics, including trend, transition, fluctuation, and outliers. Statistical attributes provide global summaries of each time series through metrics such as temporal mean, standard deviation, and initial and final values. For time series reflecting inter-cell heterogeneity, such as system-level standard deviation and Shannon entropy, we apply the same two attribute categories but tailor their specific metrics. In this context, shape attributes include amplitude level and temporal fluctuations, while statistical attributes again consist of temporal mean and standard deviation. A complete description of the designed attributes is provided in Table 1. Notably, the framework is flexible and extensible: additional attributes can be seamlessly incorporated to accommodate specific configurations or operational requirements of different LIB systems.
Time series Shape Statistics Trend Transition Fluctuation Outliers Amplitude level Initial value Final value Temporal mean Temporal standard deviation Single LIB cell ✓ ✓ ✓ ✓ ✓ ✓ ✓ ✓ System level Aggregate behavior Mean ✓ ✓ ✓ ✓ ✓ ✓ ✓ ✓ Maximum ✓ ✓ ✓ ✓ ✓ ✓ ✓ Minimum ✓ ✓ ✓ ✓ ✓ ✓ ✓ Heterogeneity Standard deviation ✓ ✓ ✓ ✓ Shannon entropy ✓ ✓ ✓ ✓
4.4 Rules for assigning attribute descriptors
To systematically assign semantic descriptors to each semantic attribute within a given temporal slice, we define a set of rule-based criteria, summarized in Table 2. These rules are grounded in established principles from standard time-series analysis literature [best2007perception, miot2020empirical, borowski2014online, bertrand2011off, cousineau2010outliers, rousseeuw2011robust], ensuring consistency and interpretability across time series. The principles for selecting hyperparameters are detailed in Supplementary Note 6 and Table S3.
Attributes Hyperparameters Assigning rules Descriptors Trend Fit the time series using a linear function . If the slope exceeds a positive threshold , the trend is classified as “increasing”; if , it is classified as “decreasing”. Otherwise, the trend is considered “stable”. Increasing; Decreasing; Stable Transition , Estimate the local trend using a sliding window of size . Within each window, perform a first-order linear fit to compute a local slope, resulting in a sequence of local slope estimates over time. A transition event is detected when the absolute difference between two successive local slopes exceeds . The temporal midpoint of the corresponding sliding window is defined as the transition time. Transition at the time stamp; Without transition Fluctuation Apply a linear fit to the data within each time slice, modeling as . Compute the detrended signal . Calculate the variance of . If this variance exceeds , classify as noticeable fluctuation. Fluctuation Remains slight; With noticeable fluctuation Outlier Outliers are detected by using the Z-score method. The absolute deviation at each time stamp is computed as , where . A time stamp is identified as an outlier if its absolute if exceeds both and a predefined threshold , where . Outlier detected at the time stamp; Without outliers Amplitude level , Compute the absolute mean value of the signal and define two thresholds, and (where ). If the absolute mean exceeds , the signal is classified as “significant”; if it lies between and , it is classified as “moderate”; otherwise, it is considered “slight”. Significant; Moderate; Slight Temporal mean - Calculate the average value of the time series across the duration of the time slice. - Temporal standard deviation - Calculate the standard deviation of the time series across the duration of the time slice. - Initial value - Record the value of the time series at the start of the time slice. - Final value - Record the value of the time series at the end of the time slice. -
4.5 Consolidation of redundant local descriptors
At this stage, each time series is represented as a sequence of semantic descriptors. To reduce redundancy while preserving descriptive quality, we perform a descriptor merging step that consolidates consecutive slices with identical semantics. Starting from the first time slice, we iteratively compare each descriptor with the next; if they are identical, the slices are merged into a single segment. For example, if the voltage of a LIB cell is consistently increasing across time slices 1 to 4, these slices are combined under the unified descriptor “increasing”.
4.6 Translating descriptors into expressions
After merging, each descriptor is paired with its corresponding time slice and target time series to form a triplet:
Descriptor: Trend – increasing
Target signal: Voltage of Cell 1
Time slice: to time stamps
This triplet is then converted into natural language using a predefined translation template (see Table S5). For the example above, the resulting expression is:
“From the to the time stamp, the voltage of Cell 1 kept increasing.”
4.7 Converting expressions into reports
After the raw LIB time-series data are transformed into textual expressions in the previous stage, these expressions still appear as a sequence of record-style entries with limited readability and weak structural coherence. To resolve this issue, we further employ LLMs to reorganize these semantic expressions into structurally rigorous, fluent, and information-complete reports. Leveraging the intrinsic capability of LLMs for text understanding, synthesis, and summarization, this conversion can be accomplished autonomously without any retraining, which fundamentally distinguishes our framework from multimodal LLM approaches for time-series analysis that typically rely on task-specific fine-tuning.
The report generation process takes two inputs: (i) the textual expressions produced in the previous stage, and (ii) an expert-guided instruction specifying the target reporting task (see Supplementary Note 7). These inputs are provided to the LLM to synthesize descriptive, multi-attribute reports corresponding to the original time-series data. This procedure is model-agnostic and can be readily implemented with any mainstream LLM through a simple API call.
To improve computational efficiency, we design the API interface to support parallel report generation across multiple single cells. Specifically, for the MIT and TJU datasets, we consider reporting tasks for 6 and 4 standalone LIB cells, respectively, and the time-series data from different cells are jointly provided as inputs, prompting the LLM to generate cell-wise reports within a single API call. For the ZJU-t dataset, a similar parallel strategy is adopted. However, due to the maximum token limitation of current LLMs, multiple independent API calls are issued to separately generate reports for groups of cells: (1) cells #1–4, (2) cells #5–8, (3) cells #9–12, and (4) cells #13–16. In addition, a system-level time-series report is generated through an independent API call. All report segments, including the four single-cell reports and the system-level report, are finally concatenated to form a complete report for the entire LIB system.
4.8 Time-series reports for downstream battery O&M applications
The generated time-series reports constitute a rich and structured textual representation of LIB operational behaviours, which can be directly integrated into mainstream LLMs for explainable battery O&M applications. Specifically, we supply the LLM with (i) task-specific human queries (see Supplementary Note 5) and (ii) the corresponding report text as prompts. Within a question–answering framework, the LLM processes these inputs to produce natural-language O&M responses.
For more specialized tasks such as fault diagnosis, expert knowledge can be further incorporated through advanced prompting strategies or RAG. This framework enables a unified, interpretable, and knowledge-enhanced interface between raw time-series data and practical battery O&M decision-making.
Conflict of interests
The authors declare that they have no conflict of interest.
Acknowledgements
J.Y. and C.Z. acknowledge support from the National Natural Science Foundation of China (Grant Nos. 62125306, 62450020, and U25B2050) and the State Key Laboratory of Industrial Control Technology, China (Grant No. ICT2025C01). Z.C. acknowledges support from a Natural Science and Engineering Research Council of Canada Discovery Grant (RGPIN-2024-06015). All authors gratefully acknowledge the members of Prof. Zhao’s group at ZJU and Prof. Furong Gao’s group at Hong Kong University of Science and Technology (HKUST) for their assistance with time-series data annotation.
Author contributions
Conceptualisation, J.Y. and C.Z.; Methodology, J.Y. and C.Z.; Software, J.Y. and C.Z.; Original data acquisition, J.Y. and C.Z.; Analysis, J.Y., C.Z., M.G. and Z.C.; Visualisation, J.Y. and C.Z.; Writing original draft, J.Y, C.Z., M.G. and Z.C.; Writing, review & editing, J.Y, C.Z., M.G. and Z.C.; Supervision, C.Z. and Z.C.